A motorboat can go 8 miles downstream on a river in 10 minutes. It takes 30 minutes for this boat to go back upstream the same 8 miles. Find the speed of the boat in still water.
Answer by scott8148(6628)

let m=motorboat and c=current

downstream __ m+c=8/(10/60) __ m+c=48 mph

upstream __ m-c=8/(30/60) __ m-c=16 mph

adding the equations __ m+c+(m-c)=48+16 __ 2m=64 __ m=32 mph